Ice has a density of 0.92 g/cm³ and seawater has a density of 1.04 g/cm³. Determine the total volume of a floating iceberg if 30 m³ is above the seawater.


1
Expert's answer
2021-05-05T07:42:03-0400

Provided that the

Density of ice is ="0.92 g\/m^3"

Density of sea water "=" "1.04 g\/cm^3"

Volume of Ice berg above water "= 30 m^3 or 3.0 \times 10^7 cm^3"

If v' is the volume of ice berg immersed in water


then as per the law of floatation = "\frac{v'}{V} = \frac{density of ice }{density of sea water}"

"=\frac{v'}{V}= \frac{0.92}{1.04}= 0.8846"



v' = 0.88361 V


Volume of the iceberg over the surface of sea water is obtained by:


V(1 - 0.8846) = "3.0 \times 10^7 cm^3"



"V = \frac{3.0 \times 10^7 cm^3}{1-0.8846} = 2.59965 \times 10^8 cm^3"

V= "259.9653 m^3"